Advertisements
Advertisements
प्रश्न
Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.
विकल्प
`186/190`
`187/190`
`188/190`
`18/(""^20C_3)`
Advertisements
उत्तर
Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive `187/190`.
Explanation:
Set of three consecutive numbers from 1 to 20 are 1, 2, 3; 2, 3, 4; 3, 4, 5; ..., 18, 19, 20.
So, the probability that the numbers are consecutive
= `18/(""^20C_3)`
= `18/((201)/(3!17!))`
= `(18*3!*17!)/(20!)`
= `(18 xx 3 xx 2 xx 17!)/(20 xx 19 xx 18 xx 17!)`
= `(3 xx 2)/(20 xx 19)`
= `3/190`
∴ P(three numbers are not consecutive) = `1 - 3/190 = 187/190`
APPEARS IN
संबंधित प्रश्न
A fair coin with 1 marked on one face and 6 on the other and a fair die are both tossed. Find the probability that the sum of numbers that turn up is (i) 3 (ii) 12
A fair coin is tossed four times, and a person win Re 1 for each head and lose Rs 1.50 for each tail that turns up.
From the sample space calculate how many different amounts of money you can have after four tosses and the probability of having each of these amounts.
In a lottery, person chooses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? [Hint: order of the numbers is not important.]
A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that all will be blue?
4 cards are drawn from a well-shuffled deck of 52 cards. What is the probability of obtaining 3 diamonds and one spade?
The number lock of a suitcase has 4 wheels, each labelled with ten digits i.e., from 0 to 9. The lock opens with a sequence of four digits with no repeats. What is the probability of a person getting the right sequence to open the suitcase?
A and B throw a pair of dice. If A throws 9, find B's chance of throwing a higher number.
Two unbiased dice are thrown. Find the probability that the total of the numbers on the dice is greater than 10.
If a letter is chosen at random from the English alphabet, find the probability that the letter is a vowel .
In a lottery, a person chooses six different numbers at random from 1 to 20, and if these six numbers match with six number already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game?
Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S = {w1, w2, w3, w4, w5, w6, w7}:
| Elementary events: | w1 | w2 | w3 | w4 | w5 | w6 | w7 |
| (iii) | 0.7 | 0.06 | 0.05 | 0.04 | 0.03 | 0.2 | 0.1 |
A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that all 10 are good
A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability thatat least one is defective
Two dice are thrown together. The probability that neither they show equal digits nor the sum of their digits is 9 will be
An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is red or yellow and numbered 1, 2, 3 or 4
An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 5, 15, 25, or 35
One mapping (function) is selected at random from all the mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is ______.
Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?
The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A)
The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine `P(A ∩ barB)`
The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(B ∩ C)
The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine Probability of exactly one of the three occurs
A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are white
If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together
If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together
While shuffling a pack of 52 playing cards, 2 are accidentally dropped. Find the probability that the missing cards to be of different colours ______.
6 boys and 6 girls sit in a row at random. The probability that all the girls sit together is ______.
The sum of probabilities of two students getting distinction in their final examinations is 1.2
The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.
| C1 Probability |
C2 Written Description |
| (a) 0.95 | (i) An incorrect assignment |
| (b) 0.02 | (ii) No chance of happening |
| (c) – 0.3 | (iii) As much chance of happening as not |
| (d) 0.5 | (iv) Very likely to happen |
| (e) 0 | (v) Very little chance of happening |
